OpenStudy (anonymous):

How do you find an equation for the line satisfying the given conditions. Through (3, 3) and parallel to 3x - 4y = 7 ??

OpenStudy (campbell_st):

write the equation of the line in point slope form y = mx + b or y = 3/4 x - 7/4 parallel lines have the same gradient... so use the gradient of this line and the point slope formula.. with the given point

OpenStudy (anonymous):

Let the equation of the line be y=mx+c m=3/4 Applying the given point and m 3=3/4*3+c c=3/4 the eqn is y=3/4x+3/4

OpenStudy (anonymous):

First isolate y: -4y=-3x+7 y=(3/4)x-7/4 The slope is 3/4. Our parallel line will have this slope as well. Point-slope form: \[y-y1=m(x-x1)\] \[y-3=\frac{3}{4}(x-3)\] slope-intercept form: \[y=mx+b\] \[y=\frac{3}{4}x+\frac{3}{4}\]